The following information has been collected from Ms. Divya the head of personnel

department of ART Limited for the year ended 31st March 2020


Number of workers at the beginning of the year -8000

Number of workers at the end of the year -9600

Number of workers left the company during the year -500

Number of workers discharged during the year -100

Number of workers replaced due to leaving and discharge -700

Additional workers employed for expansion during the year -1500


You are required to calculate the labour turnover rate by using separation method,

replacement method and flux method, further discuss each of them briefly.


1
Expert's answer
2021-05-18T06:14:04-0400

i) Separation Method



The separation method includes the necessities to separations without considering replacement made in the number of the workforce. Therefore, the labor turnover rate is determined by dividing the number of separations during a specified period and the average number of workers on the payroll during the same period. 


Labor turnover rate = "\frac{ Number of separation in a period }{ Average number of workers in the period} \times 100"


Average number of workers in the period = "\frac{(No. of labor at the beginning + No of labor at the end)}{2}" =

="\frac{8000 +9600}{2} = 8800"

Labor turnover rate= "\frac{500}{8800}\times 100= 5.68" %


ii) Replacement method


The separation method is used to measure the labor turnover ratio by dividing the realized replacement during a period by the average number of workers on the payrolls for the same time. The separation method incorporates only the number of replacements and the separations during the period. 


Labor turnover rate ="\frac{ Number of replacement in a period }{ Average number of workers in the period} \times 100"


Average number of workers in the period ="\frac{(No. of labor at the beginning + No of labor at the end)}{2}"


="\frac{8000 +9600}{2} = 8800"


Labor turnover rate= "\frac{700}{8800}\times 100= 7.95" %


iii) Flux method

The flux method considers both the number of discharged or who left and the number of workers who were replaced. The flux method combines both the Separation method and the Replacement method as follows: 




Average number of workers in the period = "\frac{(No. of labor at the beginning + No of labor at the end)}{2}"


= "\frac{8000 +9600}{2} = 8800"


Turn over Rate= "\frac{ 500 + 700 }{ 8800} \times 100= 13.64"%